OpenStudy (anonymous):

Staples sells CD's ($2.00) and boxes of computer paper ($3). If total sales were $1560 and customers bought 5 times as many CD's as boxes of computer paper, how many CD's were sold?

OpenStudy (anonymous):

let number of boxes of paper = x then number of CD's = 5x Value of x boxes of paper = x * 3 = 3x Value of 5x CD's = 5x * 2 = 10x ATQ 3x + 10x = 1560 13x = 1560 x = 1560/13 = 120 so x = 120 hence number of CD's sold = 5x = 5*120 = 600

OpenStudy (anonymous):

awesome

OpenStudy (anonymous):

u r welcome....☺